The sum of David's and Jim's ages is now 65. When David is as old as Jim is now, twice Jim's age minus David's age at that time will be 110. How old is David now?
Answer by ankor@dixie-net.com(22740)

The sum of David's and Jim's ages is now 65.
d + j = 65
let x = j-d (the difference in their ages)
When David is as old as Jim is now, twice Jim's age minus David's age at that time will be 110.
2(j+x) - (d+x)) = 110
2j + 2x - d - x = 110
2j - d + x = 110
replace x with (j-d)
2j - d +(j-d) = 110
3j-2d = 110
Use elimination, multiply the 1st equation by 2
3j - 2d = 110
2j + 2d = 130
-----------------addition eliminates d, find j
5j = 240
j = 240/5
j = 48 yrs is Jim's present age
then
48 + d = 65
d = 65 - 48
d = 17 is David's age
:
:
Check this in the statement
48 - 17 = 31 is the difference in their ages
2(48+31) - (17+31) = 110